OpenStudy (anonymous):

the position vectors of points P and Q are given by r1= 2i + 3j - k and r2= 4i - 3j + 2k. Determive PQ in terms of i,j,k and find its magnitude.

OpenStudy (anonymous):

P= <2i , 3j , -k> Q= <4i , -3j , 2k> P*Q in terms of i, j, k is pretty simple. Just multiply the two i terms to get a new i term, multiply the j terms to get a new one, multiply the k terms to get a new k, then put them in order: New Vector = <(2i)(4i) , (3j)(-3j) , (-k)(2k)> = <8i , -6j , -2k> In order to get the MAGNITUDE of P*Q, what you do is multiply like terms and add them together: (2i)(4i) + (3j)(-3j) + (-k)(2k) = (8) + (-6) + (-2) = 8-6-2 = 0 I think that should be right.

OpenStudy (anonymous):

3(-3) is not equal to 6

OpenStudy (anonymous):

You're right. It's equal to (-6), like I said.

OpenStudy (anonymous):

If you are adding a negative number, it's the same as subtracting the positive number.